# Interaction-induced phase transitions and critical phases in nonreciprocal non-Hermitian quasicrystals

### v2

@@ III.2 Asymmetric hopping model (AHM)
In Fig. 5, we present the spectra of $\hat{H}$ [Eq. (1)] and the IPRs of its eigenstates for typical cases with $U\neq 0$. Qualitatively, we see that the spectral and state features of the UHM are maintained even with nonzero $J_{{\rm L}}$ and $J_{{\rm R}}$. Specially, we find that when $J_{{\rm L}}<V$, the spectrum of $\hat{H}$ is real and all its eigenstates are localized for $U\neq 0$. Since the onsite interaction only promotes localization, and localized bosons are controlled by the quasiperiodic potential, the spectral properties observed in Figs. 5(a), 5(c) and 5(e) are expected. At large $U$, the spectrum also divides into two portions gapped along ${\rm Re}E$ as that of UHM. The states with higher energies are bosonic doublons localized at different lattice sites. For $J_{{\rm L}}>V$, large amounts of eigenenergies become complex. With the rise of $U$, some energies start to leave the main body of spectrum, composing a loop at the highest energy of the noninteracting AHM, and then pinned to ${\rm Re}E$. Similar to UHM, the states with real energies due to a nonzero $U$ are spatially localized. We should then have an interaction-induced critical phase in the AHM with coexisting extended and localized eigenstates, which disappears when $U=0$. Further increasing $U$ yields $L$ localized states with real energies. Finally, the spectrum becomes two disconnected sets containing extended states with complex energies and localized bosonic doublons with real energies, respectively. These observations suggest that generic features of the UHM regarding spectral, localization and topological transitions should be carried over to the AHM with both $J_{{\rm L}},J_{{\rm R}}\neq 0$.
@@
Figure 6: Spectral and localization transitions of the AHM vs $(J_{{\rm L}},U)$ under the PBC. Other system parameters are $(J_{{\rm R}},V)=(1,1.5)$ and the length of lattice is $L=89$ for all panels. (a) Ratio of eigenenergies with nonzero imaginary parts $\rho_{{\rm Im}}$ [Eq. (3)]. (b) The maximum of IPRs ${\rm IPR}_{\max}$ [Eq. (7)]. (c) The minimum of IPRs ${\rm IPR}_{\min}$ [Eq. (8)]. (d) The smoking-gun function $\zeta$ [Eq. (9)] of the critical phase.
@@
To gain a more comprehensive understanding, we evaluate the functions $\rho_{{\rm Im}}$ [Eq. (3)], ${\rm IPR}_{\max}$ [Eq. (7)], ${\rm IPR}_{\min}$ [Eq. (8)] and $\zeta$ [Eq. (9)] in a typical domain of the parameter space $(J_{{\rm L}},U)$ for the AHM, with results presented in Fig. 6. In Fig. 6(a), we realize that the real-to-complex spectral transition of the noninteracting model is essentially not affected by interactions, in the sense that the bulk of eigenenergies turn into complex at any $U\neq 0$ across the transition point $J_{{\rm L}}=V=1.5$ of the $\hat{H}_{0}\equiv\hat{H}(U=0)$ [Eq. (1)]. However, this is not the case for the delocalization transition. In Fig. 6(c), we find that extended states start to appear across the spectral transition at $J_{{\rm L}}=V=1.5$, indicating the onset of delocalization transitions. Nevertheless, the system could enter the extended phase only at a $J_{{\rm L}}$ higher than the noninteracting critical point $V=1.5$ for any $U\neq 0$, as shown by the transition boundary of the ${\rm IPR}_{\max}$ in Fig. 6(b). Therefore, onsite interactions could drive the ending point of delocalization transition of the AHM towards a stronger asymmetric hopping regime. Similar situations have been encountered before in the last subsection, highlighting the generality of local-interaction effects on NHQCs. Altogether, this means that the joint critical point of the spectral and localization transitions of the noninteracting AHM [Eq. (1)] would be split and extended into a critical phase, carrying both extended states of complex energies and localized states of real energies, with two such examples shown in Figs. 5(d) and 5(f). The presence of such a critical phase is further affirmed by the smoking-gun function $\zeta$ in Fig. 6(d), which shows finite values only in the region between $J_{{\rm L}}=V=1.5$ [the left boundary of $\rho_{{\rm Im}}$ and ${\rm IPR}_{\min}$] and the right boundary of ${\rm IPR}_{\max}$, i.e., a region sandwiched between the extended and localized phases.
@@
Finally, we check the winding numbers $(\omega_{1},\omega_{2})$ in Eq. (10) for the AHM, with results presented in Fig. 7. In the localized phase (with $1.50\leq J_{{\rm L}}<$ 1.5 at any $U$), the spectrum of AHM is real with no loops on the complex plane, and we find $\omega_{1}=\omega_{2}=0$. In the critical phase, complex energy loops start to appear in the spectrum, and we find $\omega_{1}>0$ yet $\omega_{2}=0$. The change of $\omega_{1}$ from zero to a finite integer thus provides the topological signature for the transition from localized to critical phases. In the extended phase, we find both $\omega_{1}>0$ and $\omega_{2}>0$. Therefore, the jump of $\omega_{2}$ from zero to a finite integer offers the topological signal for the transition from critical to extended phases. The combination of information provided by $\omega_{1}$ and $\omega_{2}$ then allows us to have a complete topological characterization of the spectral and localization transitions in the AHM, which is also consistent with the description based on other quantities in Fig. 6. Note in passing that the signs of winding numbers $(\omega_{1},\omega_{2})$ here are different from those of the UHM in the last subsection. The underlying reason is that in the extended and critical phases, the skin-localization directions of bulk states in the UHM and AHM are opposite. Under the OBC, the eigenstates are localized at the right edge in the extended and critical phases of the UHM, while they are localized at the left edge in these phases of the AHM for our considered parameter space.
@@
To sum up, in two typical nonreciprocal AAH quasicrystals, the cooperation among disorder, interaction and non-Hermitian effects could indeed create unique types of spectral, localization and topological transitions. A key role of the onsite interaction is to dissolve the joint critical point of single-particle spectral and localization transitions and transform it into a critical phase with both extended and localized states. In the strong-interaction regime, this critical phase is expected to hold two disconnected clusters of states. One of them contains extended two-body states, whose spectrum looks similar to the associated noninteracting system. The other one consists of bounded boson pairs forming doublons. We explore the spectral, localization and topological physics of these bosonic doublons in the next section.
@@ V Conclusion
In this study, we uncovered rich spectral and localization phenomena driven by the interplay among nonreciprocal hopping, correlated disorder and interparticle interactions in a representative class of non-Hermitian quasicrystal. With a focus on two bosons subject to Hubbard onsite interactions in AAH-type lattices, we revealed that interactions could enhance localization, specially pushing the threshold of localization transition towards stronger hopping regions. Moreover, interactions could give rise to a critical phase, in which extended two-body states and bounded boson pairs are coexistent. Meanwhile, the transitions between extended, critical and localized phases in our system are topological in nature and characterized by a pair of integer quantized winding numbers. Finally, we analyzed the doublon states under strong interactions and unveiled their non-Hermitian skin effects, whose direction could be further controlled flexibly via tuning the asymmetric hopping parameters.

### v3

@@ III.2 Asymmetric hopping model (AHM)
In Fig. 5, we present the spectra of $\hat{H}$ [Eq. (1)] and the IPRs of its eigenstates for typical cases with $U\neq 0$. Qualitatively, we see that the spectral and state features of the UHM are maintained even with nonzero $J_{{\rm L}}$ and $J_{{\rm R}}$. Specially, we find that when $J_{{\rm L}}<V$, the spectrum of $\hat{H}$ is real and all its eigenstates are localized for $U\neq 0$. Since the onsite interaction only promotes localization, and localized bosons are controlled by the quasiperiodic potential, the spectral properties observed in Figs. 5(a), 5(c) and 5(e) are expected. At large $U$, the spectrum also divides into two portions gapped along ${\rm Re}E$ as that of UHM. The states with higher energies are bosonic doublons localized at different lattice sites. For $J_{{\rm L}}>V$, large amounts of eigenenergies become complex. With the rise of $U$, some energies start to leave the main body of spectrum, composing a loop at the highest energy of the noninteracting AHM, and then pinned to ${\rm Re}E$. Similar to UHM, the states with real energies due to a nonzero $U$ are spatially localized. We should then have an interaction-induced intermediate mobility edge phase in the AHM with coexisting extended and localized eigenstates, which disappears when $U=0$. Further increasing $U$ yields $L$ localized states with real energies. Finally, the spectrum becomes two disconnected sets containing extended states with complex energies and localized bosonic doublons with real energies, respectively. These observations suggest that generic features of the UHM regarding spectral, localization and topological transitions should be carried over to the AHM with both $J_{{\rm L}},J_{{\rm R}}\neq 0$.
@@
Figure 6: Spectral and localization transitions of the AHM vs $(J_{{\rm L}},U)$ under the PBC. Other system parameters are $(J_{{\rm R}},V)=(1,1.5)$ and the length of lattice is $L=89$ for all panels. (a) Ratio of eigenenergies with nonzero imaginary parts $\rho_{{\rm Im}}$ [Eq. (3)]. (b) The maximum of IPRs ${\rm IPR}_{\max}$ [Eq. (7)]. (c) The minimum of IPRs ${\rm IPR}_{\min}$ [Eq. (8)]. (d) The smoking-gun function $\zeta$ [Eq. (9)] of the intermediate mobility edge phase.
@@
To gain a more comprehensive understanding, we evaluate the functions $\rho_{{\rm Im}}$ [Eq. (3)], ${\rm IPR}_{\max}$ [Eq. (7)], ${\rm IPR}_{\min}$ [Eq. (8)] and $\zeta$ [Eq. (9)] in a typical domain of the parameter space $(J_{{\rm L}},U)$ for the AHM, with results presented in Fig. 6. In Fig. 6(a), we realize that the real-to-complex spectral transition of the noninteracting model is essentially not affected by interactions, in the sense that the bulk of eigenenergies turn into complex at any $U\neq 0$ across the transition point $J_{{\rm L}}=V=1.5$ of the $\hat{H}_{0}\equiv\hat{H}(U=0)$ [Eq. (1)]. However, this is not the case for the delocalization transition. In Fig. 6(c), we find that extended states start to appear across the spectral transition at $J_{{\rm L}}=V=1.5$, indicating the onset of delocalization transitions. Nevertheless, the system could enter the extended phase only at a $J_{{\rm L}}$ higher than the noninteracting critical point $V=1.5$ for any $U\neq 0$, as shown by the transition boundary of the ${\rm IPR}_{\max}$ in Fig. 6(b). Therefore, onsite interactions could drive the ending point of delocalization transition of the AHM towards a stronger asymmetric hopping regime. Similar situations have been encountered before in the last subsection, highlighting the generality of local-interaction effects on NHQCs. Altogether, this means that the joint critical point of the spectral and localization transitions of the noninteracting AHM [Eq. (1)] would be split and extended into an intermediate mobility edge phase, carrying both extended states of complex energies and localized states of real energies, with two such examples shown in Figs. 5(d) and 5(f). The presence of such an intermediate phase is further affirmed by the smoking-gun function $\zeta$ in Fig. 6(d), which shows finite values only in the region between $J_{{\rm L}}=V=1.5$ [the left boundary of $\rho_{{\rm Im}}$ and ${\rm IPR}_{\min}$] and the right boundary of ${\rm IPR}_{\max}$, i.e., a region sandwiched between the extended and localized phases.
@@
Finally, we check the winding numbers $(\omega_{1},\omega_{2})$ in Eq. (10) for the AHM, with results presented in Fig. 7. In the localized phase (with $1.50\leq J_{{\rm L}}<$ 1.5 at any $U$), the spectrum of AHM is real with no loops on the complex plane, and we find $\omega_{1}=\omega_{2}=0$. In the intermediate phase, complex energy loops start to appear in the spectrum, and we find $\omega_{1}>0$ yet $\omega_{2}=0$. The change of $\omega_{1}$ from zero to a finite integer thus provides the topological signature for the transition from localized to intermediate phases. In the extended phase, we find both $\omega_{1}>0$ and $\omega_{2}>0$. Therefore, the jump of $\omega_{2}$ from zero to a finite integer offers the topological signal for the transition from intermediate to extended phases. The combination of information provided by $\omega_{1}$ and $\omega_{2}$ then allows us to have a complete topological characterization of the spectral and localization transitions in the AHM, which is also consistent with the description based on other quantities in Fig. 6. Note in passing that the signs of winding numbers $(\omega_{1},\omega_{2})$ here are different from those of the UHM in the last subsection. The underlying reason is that in the extended and intermediate phases, the skin-localization directions of bulk states in the UHM and AHM are opposite. Under the OBC, the eigenstates are localized at the right edge in the extended and intermediate phases of the UHM, while they are localized at the left edge in these phases of the AHM for our considered parameter space.
To sum up, in two typical nonreciprocal AAH quasicrystals, the cooperation among disorder, interaction and non-Hermitian effects could indeed create unique types of spectral, localization and topological transitions. A key role of the onsite interaction is to dissolve the joint critical point of single-particle spectral and localization transitions and transform it into an intermediate mobility edge phase with both extended and localized states. In the strong-interaction regime, this intermediate phase is expected to hold two disconnected clusters of states. One of them contains extended two-body states, whose spectrum looks similar to the associated noninteracting system. The other one consists of bounded boson pairs forming doublons. We explore the spectral, localization and topological physics of these bosonic doublons in the next section.
@@ V Conclusion
In this study, we uncovered rich spectral and localization phenomena driven by the interplay among nonreciprocal hopping, correlated disorder and interparticle interactions in a representative class of non-Hermitian quasicrystal. With a focus on two bosons subject to Hubbard onsite interactions in AAH-type lattices, we revealed that interactions could enhance localization, specially pushing the threshold of localization transition towards stronger hopping regions. Moreover, interactions could give rise to an intermediate mobility edge phase, in which extended two-body states and bounded boson pairs are coexistent. Meanwhile, the transitions between extended, intermediate and localized phases in our system are topological in nature and characterized by a pair of integer quantized winding numbers. Finally, we analyzed the doublon states under strong interactions and unveiled their non-Hermitian skin effects, whose direction could be further controlled flexibly via tuning the asymmetric hopping parameters.
With more than two interacting bosons, there could also be interaction-induced bound states. Intermediate mobility edge phases are thus expected to emerge with the increase of disorder or interaction strengths. For example, our preliminary calculations with three interacting bosons yield an intermediate phase, in which localized triplons (three-boson bound states) coexist with extended states. They are also subject to NHSEs under the OBC. In the strong interaction regime, as the effective hopping amplitude of triplons is proportional to $J^{2}_{\rm L,R}/U$, they will be first localized with the increase of disorder, leading to the localization transition at weaker disorder strengths compared to the noninteracting case. Therefore, although we focus on the case of two interacting bosons in this study, the physical mechanisms underlying the interaction-induced intermediate phases and the NHSEs of multiparticle bound states we revealed are expected to hold in the cases of more than two particles.